# R&D Networks under Heterogeneous Firm Productivities

### v1

@@ Appendix B Additional Simulation Results
Note: This figure isolates the effect of changing the productivity composition while holding the network fixed. We consider $n=10$ firms partitioned into two fully connected components (two cliques of size 55), mirroring the connectivity pattern of the Positive Assortative (PA) network at $\rho=\tfrac{1}{2}$. Starting from $\rho=0$ (all firms have low productivity $\theta$), we increase $\rho$ in increments of $0.1$ by upgrading one firm at a time from $\theta$ to 11 (high productivity). At each step, we record the change in profit for the transitioning firm (the firm whose productivity is upgraded in that step), holding all other firms’ productivities and the network links fixed. The x-axis reports $\rho$ (the share of high-productivity firms), and the y-axis reports the corresponding profit change of the transitioning firm. Curves are shown for three values of the low productivity parameter, $\theta\in\{0.1,0.5,0.9\}$. Two patterns emerge. First, the profit gain from upgrading is larger when $\theta$ is smaller, consistent with a larger productivity jump from $\theta$ to $1$. Second, the profit gain is generally increasing in $\rho$, except around $\rho=\tfrac{1}{2}$, where the fixed network aligns with exact type clustering (all high types concentrated within one clique and all low types within the other); upgrading a low-type firm within the low-type clique yields a smaller marginal benefit than upgrading the remaining low-type firm in the high-type clique. In all cases the profit change remains positive. Finally, note that this fixed-network exercise is not, in general, a stability analysis: the imposed two-clique network need not be pairwise stable for the corresponding productivity composition, except at $\rho=\tfrac{1}{2}$ (and sufficiently low $\theta$).

### v2

@@ Appendix B Additional Simulation Results
Note: This figure isolates the effect of changing the productivity composition while holding the network fixed. We consider $n=10$ firms partitioned into two fully connected components (two cliques of size 55), mirroring the connectivity pattern of the Positive Assortative (PA) network at $\rho=\tfrac{1}{2}$. Starting from $\rho=0$ (all firms have low productivity $\theta$), we increase $\rho$ in increments of $0.1$ by upgrading one firm at a time from $\theta$ to 11 (high productivity). At each step, we record the change in profit for the transitioning firm (the firm whose productivity is upgraded in that step), holding all other firms’ productivities and the network links fixed. The x-axis reports $\rho$ (the share of high-productivity firms), and the y-axis reports the corresponding profit change of the transitioning firm. Curves are shown for three values of the low productivity parameter, $\theta\in\{0.1,0.5,0.9\}$. Two patterns emerge. First, the profit gain from upgrading is larger when $\theta$ is smaller, consistent with a larger productivity jump from $\theta$ to 11. Second, the profit gain is generally increasing in $\rho$, except around $\rho=\tfrac{1}{2}$, where the fixed network aligns with exact type clustering (all high types concentrated within one clique and all low types within the other); upgrading a low-type firm within the low-type clique yields a smaller marginal benefit than upgrading the remaining low-type firm in the high-type clique. In all cases the profit change remains positive. Finally, note that this fixed-network exercise is not, in general, a stability analysis: the imposed two-clique network need not be pairwise stable for the corresponding productivity composition, except at $\rho=\tfrac{1}{2}$ (and sufficiently low $\theta$).